OpenStudy (anonymous):

Solve for x: \left( \frac{x+63}{x-127} \right) ^2 - 175 \left( \frac{x+63}{x-127} \right) + 5934 = 0 The smaller solution is _____. The larger solution is ____. Please explain how to solve!

OpenStudy (anonymous):

\[\left( \frac{x+63}{x-127} \right) ^2 - 175 \left( \frac{x+63}{x-127} \right) + 5934 = 0 \]

OpenStudy (homeworksucks):

\[\left( \frac{ x+63 }{ x-127 }\right)^{2}-175(\frac{ x+63 }{ x-127})+5934=0\]

OpenStudy (homeworksucks):

Beat me to it

OpenStudy (anonymous):

lol

OpenStudy (homeworksucks):

Umm, for this question you're going to have to simplify each bracket

OpenStudy (ranga):

Make the substitution y = (x + 63) / (x - 27). It will be a simple quadratic equation in y. Solve for y. And then put y = (x + 63) / (x - 27) back and solve for x.

OpenStudy (anonymous):

let me know if you still need this solved

OpenStudy (anonymous):

Yes. I can't seem to get it right.

OpenStudy (anonymous):

Find the reference angle for csc(theta) = -2.7885